# Agile Decision-Making and Safety-Critical Motion Planning for Emergency Autonomous Vehicles

### v3

@@ V-B Benchmark and Metrics
    MOBIL [38] MOBIL is a well-known lane change strategy designed to improve overall traffic flow efficiency, the speed benefits for the ego vehicle, and enhance safety. It achieves these goals by carefully managing lane changes to minimize the braking impact on surrounding vehicles, thereby maintaining steady traffic conditions.
@@
$(m)$ $(m)$ $(m)$ $(m/s)$ $(m/s)$ $(m)$ $(m)$ $(m/s^{2})$ $(m/s^{2})$ $(m/s^{3})$
@@
V-C Statistical Results
We compare our methods against the baselines discussed in Section V-B. The statistical results for the evaluated metrics are summarized in Table III.
[…]
Efficiency: Regarding efficiency, IDEAM reached an average velocity of $12.44$ $m/s$, outperforming all other methods. It maintained a speed advantage of approximately $0.53$ $m/s$ to $1.11$ $m/s$ over the others. Additionally, IDEAM achieved the highest maximum velocity of $21.32$ $m/s$, further demonstrating its capability to optimize vehicle performance across different conditions. This superior speed and acceleration highlight IDEAM’s strong potential in emergency scenarios, where minimizing response time and enhancing operational efficiency are crucial.
@@ V-C Statistical Results
Comfort: Comfort is evaluated by analyzing vehicle dynamics, including maximum acceleration, average acceleration, and average jerk. The table shows that all methods have the same maximum acceleration of $3.00$ $m/s^{2}$. In terms of average acceleration, the SO-DM ($LAS=60$) and DRB-FSM methods performed best, both with the lowest average acceleration of $0.87$ $m/s^{2}$. In contrast, IDEAM had a higher average acceleration of $1.40$ $m/s^{2}$. Regarding average jerk, which measures the smoothness of acceleration, the DRB-FSM method performed the best with a value of $0.08$ $m/s^{3}$, offering the smoothest driving experience. IDEAM had a higher average jerk of $0.22$ $m/s^{3}$, which indicates a slight trade-off in comfort compared to SO-DM, DRB-FSM, and MOBIL. These results suggest that while IDEAM maintains strong performance in efficiency and progress, it shows a modest difference in ride smoothness.
[…]
@@
V-D Case Study
In this section, we randomly choose an interesting traffic scenario to illustrate the effectiveness of the proposed method further. By analyzing the vehicle behavior and decision-making process in this scenario, we can gain a more intuitive understanding of how the method performs in dense traffic situations. We selected three key frames at $t=13.7s$, $t=17.2s$, and $t=30.1s$, as shown in Fig  9. At $t=13.7s$, the IDEAM method, after performing LSGM to search for the desired group and then conducting lane-probing, ultimately led the ego vehicle to do a left lane change to achieve greater speed benefits. In contrast, the No-Probing IDEAM did not actively explore spatial benefits, resulting in a later lane change compared to IDEAM. The other three baselines did not perform a lane change at $t=13.7s$. SO-DM ($LAS=60$) and DRB-FSM executed a left lane change at the beginning because the leading vehicle in the left lane had an initial speed of $11.3$ $m/s$ and maintained a speed advantage over the leading vehicles in the adjacent lanes for the next five time steps. Therefore, SO-DM ($LAS=60$) and DRB-FSM chose the left lane early on to capitalize on the speed benefits, while SO-DM ($LAS=30$) initially focused on the spatial benefits of its current lane before later opting for the right lane’s efficiency gain. Given the advantage from the earlier lane change at $t=13.7s$ seconds, at $t=17.2s$ seconds, the IDEAM method guided the ego vehicle to change lanes to the left, while the No-Probing IDEAM had just moved the vehicle to the middle lane. The remaining three baselines did not change lanes between $t=13.7s$ seconds and $t=17.2s$ seconds, which reflects their lack of proactive exploration or overly conservative decision-making, as they considered the small distance between the rear vehicles in the adjacent lanes and the ego vehicle. The IDEAM method enabled the ego vehicle to complete its fourth lane change by $t=30.1s$ seconds, achieving the greatest progress among all methods. Additionally, this demonstrates that SO-DM ($LAS=30$) is more suited for dense traffic compared to SO-DM ($LAS=60$), as it better leverages spatial benefits in such scenarios. The final lane change in the IDEAM method can be seen as an overtaking maneuver of the yellow vehicle in the middle lane, while all other methods still lag in overtaking this vehicle. Additionally, IDEAM completed $5$ lane changes in total, while SO-DM ($LAS=30$) made $3$ lane changes, SO-DM ($LAS=60$) and DRB-FSM made $2$ lane changes each, and MOBIL executed $3$ lane changes. No-Probing IDEAM also made $5$ lane changes, but its performance was less efficient than IDEAM, further demonstrating the advantages of proactive lane-probing in dynamic traffic environments. Fig. 8 presents the acceleration, lateral velocity, and longitudinal velocity of all methods in this case.
VI Conclusion

### v4

@@ VI-B Benchmark and Metrics
    MOBIL [24] MOBIL is a well-known traditional lane change strategy designed to improve overall traffic flow efficiency, the speed benefits for the ego vehicle, and enhance safety. It achieves these goals by carefully managing lane changes to minimize the braking impact on surrounding vehicles, thereby maintaining steady traffic conditions. This method is primarily focused on optimizing traffic flow in typical driving scenarios, distinguishing it from methods designed specifically for high-priority situations like emergency autonomous vehicles.
@@
$(\mathrm{m})$ $(\mathrm{m})$ $(\mathrm{m})$ $(\mathrm{m/s})$ $(\mathrm{m/s})$ $(\mathrm{m})$ $(\mathrm{m})$ $(\mathrm{m/s}^{2})$ $(\mathrm{m/s}^{2})$ $(\mathrm{m/s}^{3})$
@@
VI-C Statistical Results
We compare our methods against the baselines discussed in Section VI-B. The statistical results for the evaluated metrics are summarized in Table III.
Progress: In terms of progress, the IDEAM performs best among all evaluation metrics. In the average $20~{}\mathrm{s}$ and $40~{}\mathrm{s}$ progress measurements, IDEAM achieved $265.08$ meters and $525.19$ meters, respectively, significantly ahead of other methods, indicating its continued superiority on long-time scales. Compared with No-probing IDEAM, IDEAM demonstrated a clear advantage, with a difference of $7.82$ meters at $20~{}\mathrm{s}$ and $26.22$ meters at $40~{}\mathrm{s}$. This difference highlights the benefits of incorporating lane probing in the motion planner design, allowing IDEAM to make more informed decisions and optimize vehicle trajectory more effectively. In comparison to SO-DM, DRB-FSM, and MOBIL, IDEAM opened up an even larger gap. At $20~{}\mathrm{s}$, IDEAM extended its lead over these methods by a range of approximately $11$ to $14$ meters. By $40~{}\mathrm{s}$, this gap widened further, with IDEAM ahead by about $50$ to $60$ meters. These results highlight the effectiveness of the IDEAM method in maintaining superior progress over both short and long time scales compared to the other baseline methods. Furthermore, the maximum progress of the IDEAM is $619.60$ meters, which is also higher than all baselines.
Efficiency: Regarding efficiency, IDEAM reached an average velocity of $12.44$ ${}\mathrm{m/s}$, outperforming all other methods. It maintained a speed advantage of approximately $0.53~{}\mathrm{m/s}$ to $1.11~{}\mathrm{m/s}$ over the others. Additionally, IDEAM achieved the highest maximum velocity of $21.32~{}\mathrm{m/s}$, further demonstrating its capability to optimize vehicle performance across different conditions. This superior speed and acceleration highlight IDEAM’s strong potential in emergency scenarios, where minimizing response time and enhancing operational efficiency are crucial.
@@ VI-C Statistical Results
Comfort: Comfort is evaluated by analyzing vehicle dynamics, including maximum acceleration, average acceleration, and average jerk. The table shows that all methods have the same maximum acceleration of $3.00~{}\mathrm{m/s}^{2}$. In terms of average acceleration, the SO-DM ($LAS=60$) and DRB-FSM methods performed best, both with the lowest average acceleration of $0.87~{}\mathrm{m/s}^{2}$. In contrast, IDEAM had a higher average acceleration of $1.40~{}\mathrm{m/s}^{2}$. Regarding average jerk, which measures the smoothness of acceleration, the DRB-FSM method performed the best with a value of $0.08~{}\mathrm{m/s}^{3}$, offering the smoothest driving experience. IDEAM had a higher average jerk of $0.22$ ${}\mathrm{m/s}^{3}$, which indicates a slight trade-off in comfort compared to SO-DM, DRB-FSM, and MOBIL. These results suggest that while IDEAM maintains strong performance in efficiency and progress, it shows a modest difference in ride smoothness.
Fig. 8 shows the boxplots of the $20~{}\mathrm{s}$ and $40~{}\mathrm{s}$ progress, velocity, and $S_{o}$ (safety) metrics of the proposed IDEAM method compared against other baseline methods. Although IDEAM exhibits slightly smaller $S_{o}$ values, these values remain within acceptable safety margins, ensuring that the system maintains a sufficient level of safety while delivering significant improvements in efficiency and progress.
@@
Computation Time: As shown in Fig. 7, the LSGM algorithm achieves an average execution time of $0.34~{}\mathrm{ms}$ and a maximum of $2.01~{}\mathrm{ms}$, which includes the runtime of the DFS module. These values are based on statistics collected from each frame across several scenarios. The DFS module itself incurs negligible computational cost, with an average close to zero and a maximum of $1.17~{}\mathrm{ms}$. The solver exhibits an average execution time of $5.97~{}\mathrm{ms}$ and a maximum of $21.03~{}\mathrm{ms}$.
Figure 7: Computation Time of the solver, LSGM algorithm, and the C-DFS algorithm.
VI-D Case Study
In this section, two types of scenarios, namely evaluation and emergency scenarios, are presented.
Evaluation Scenario We randomly select a scenario to intuitively demonstrate the effectiveness of the proposed method. Specifically, we selected three key frames at $t=13.7~{}\mathrm{s}$, $t=17.2~{}\mathrm{s}$, and $t=30.1~{}\mathrm{s}$, as shown in Fig. 10. At $t=13.7~{}\mathrm{s}$, after lane-probing, IDEAM guided the ego vehicle to make an early left lane change for speed gain. In contrast, No-Probing IDEAM and MOBIL delayed its lane change due to the lack of spatial exploration. SO-DM ($LAS=60$) and DRB-FSM also performed early left lane changes, benefiting from a faster leading vehicle ($11.3~{}\mathrm{m/s}$), whereas SO-DM ($LAS=30$) prioritized spatial gaps and shifted to the right lane later. By $t=17.2~{}\mathrm{s}$, IDEAM had already completed another left lane change, while No-Probing IDEAM had just reached the middle lane, and other baselines remained unchanged due to conservative or passive decisions. At $t=30.1~{}\mathrm{s}$, IDEAM completed another lane change, overtaking the yellow vehicle in the middle lane and achieving the greatest progress. Conversely, MOBIL and other baselines lagged without effectively completing the overtaking maneuver. Notably, SO-DM ($LAS=30$) outperformed SO-DM ($LAS=60$) by making better use of spatial benefits. Overall, IDEAM completed five lane changes, while SO-DM ($LAS=30$), SO-DM ($LAS=60$), DRB-FSM, and MOBIL performed three, two, two, and three lane changes, respectively. Although No-Probing IDEAM also executed five lane changes, its performance was less efficient, demonstrating the advantage of proactive lane-probing. All methods’ acceleration, lateral velocity, and longitudinal velocity profiles are shown in Fig. 9.
Emergency Scenario As illustrated in Fig. 11, two emergency scenarios are presented where the ego vehicle encounters sudden deceleration of a leading vehicle. In the first scenario, the leading vehicle gradually slows down, prompting the ego vehicle to reduce its speed while actively exploring more favorable lanes. At $t=13.2~{}\mathrm{s}$, it performs a left lane change, safely avoids the slowing vehicle, and subsequently overtakes it to regain speed advantages. In the second scenario, the leading vehicle undergoes a sudden and sharp deceleration, forcing the ego vehicle to promptly slow down to avoid a potential collision. Despite the urgency, the ego vehicle continues to search for safer and more efficient lanes and successfully executes a left lane change at $t=16.0~{}\mathrm{s}$ to bypass the decelerating vehicle. It then overtakes the lead vehicle and restores its driving efficiency.
VII Discussion
The IDEAM framework has several limitations that need to be addressed. These limitations lead us to consider future work in improving the system’s ability. Firstly, it does not fully account for the various sources of uncertainty, such as environmental factors, sensor noise, and interaction uncertainties. Environmental uncertainties, like road conditions, may significantly influence vehicle behavior modeling in motion planning. While methods like Gaussian process regression may help model these uncertainties, improving the vehicle dynamics simulation’s accuracy, further work is needed to refine this integration.
Additionally, interaction uncertainty, especially the variability in reaction times, plays a crucial role in decision-making. As highlighted in [59], human reaction times tend to average around $1.3~{}\mathrm{s}$, while autonomous vehicles respond in approximately $100$ milliseconds. This difference may be incorporated into the DHOCBF and DCBF formulations to improve real-world adaptability. Furthermore, sensor noise and perception errors affect tracking and prediction, and could potentially be mitigated by integrating techniques such as Kalman filtering into the planner to enhance the planner’s performance in uncertain environments. Future work could aim to refine these approaches to create a more robust and adaptive system.
Furthermore, the treatment of interacting agents can be further developed. Game-theoretic approaches could be employed to model and optimize interactions among agents, enhancing the system’s adaptability in high-interaction environments.
VIII Conclusion
@@
(a) Snapshots from the first emergency scenario.
(b) Snapshots from the second emergency scenario.
Figure 11: Snapshots from two emergency scenarios at certain timestamps.